Hello Tutor,
I am in need of some help if you can help me out I would appreciate it.
Solve by completing the square:
1.) 4x^2 + x - 3 = 0 I got -3/4 , 1 is this correct?
2.) x^2 = 5x + 2 I think this is 5 +- sqrt of 33 / 2 is this correct?
3.) 4x^2 + 2x - 3 = 0 I got -1+- sqrt of 13 / 8 is this correct?
I am not sure I am solving these questions of completing the squares correctly if you could tell me if I have them correct I would appreciate it. Thanks in advance.

Answer by edjones(8007)

4x^2+x-3=0
x^2+x/4-3/4=0 divide both sides by 4.
x^2+x/4 =3/4
x^2+x/4+1/64=3/4+1/64 divide the 2nd term coefficient by 2 and square it and add it to both sides.
(x+1/8)^2=49/64 factor
x+1/8=+-7/8 square root of each side.
x=-1/8+-7/8
x=3/4, x=-1
you shouldn't have any problem doing the other 2.
Ed